OpenStudy (anonymous):

REACH THE IMPOSSIBLE DREAM In the situation in which somehow, in some fashion, by some trick - this 2 different series (see below) had a \[\color{red}{\bf FINITE\quad Total \quad Sum}\], what would such Total Sums be ? First series \[\sum_{k=1}^\infty \frac{1}{k}\] Second series \[\sum_{k=1}^\infty (-1)^{k+1}\] What would the sums be ?

OpenStudy (anonymous):

Solutions Exists.

OpenStudy (anonymous):

But this is infinite @sauravshakya

OpenStudy (anonymous):

id like to know what is solution...so watching :)

OpenStudy (anonymous):

ME too.

OpenStudy (anonymous):

Well there was a MAN from your part of the world... VERY famous.... mystical mathematician... numbers were his personal friends 1) Who was HE ? 2) How does he help you in that ?

OpenStudy (anonymous):

CALCULATOR

OpenStudy (anonymous):

@estudier

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Srinivasa_Ramanujan

OpenStudy (anonymous):

Are u sure the question is correct....... i think it is 1/k instead of 1/n

OpenStudy (anonymous):

Cesaro/Grandi

OpenStudy (anonymous):

Abel

OpenStudy (anonymous):

You get nothing @estudier . You are the destructor of PLEASURES

OpenStudy (anonymous):

THE SECOND ONE IS 0

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

I think for first one it is greater than 2

OpenStudy (anonymous):

Divergent series upset the calculs addicts....

OpenStudy (anonymous):

\[ Lao\,\, Tzu\,\,\,says\,\,\, \text{There is no such thing as divergent series} \]

OpenStudy (anonymous):

No, that was Sum Tzu..

OpenStudy (anonymous):

Rubbish estudier, tat guy was only a stupid general

OpenStudy (anonymous):

Um, you didn't get my little joke....(sum not sun)

OpenStudy (anonymous):

Yes I got it - but the mind was faster then the hand.

OpenStudy (anonymous):

WELL, if @Mikael THE ONE WHO KNOWS EVERYTHING HAS ALREADY SOLVED THEN WHY SHOULD I DO IT......

OpenStudy (anonymous):

Look Abel summation or Cesaro summation - AND PLEASE SOLVE IT here. PLease!

OpenStudy (anonymous):

Grandi series

OpenStudy (anonymous):

The real understanding comes from Abel summation !

OpenStudy (anonymous):

Cesaro

OpenStudy (anonymous):

I thought everyone knew that one.....

OpenStudy (anonymous):

im back... lol

OpenStudy (anonymous):

@mukushla is brilliant in solving problem.......... @Mikael

OpenStudy (anonymous):

One of best talent I have seen.

OpenStudy (anonymous):

But you didn't look up Cesaro or Abel summation

OpenStudy (anonymous):

1+1+2+5+14+42+...= 1/2 - i sqrt3/2

OpenStudy (anonymous):

@sauravshakya @mukushla this sneaky French guy (cannot say his name) is stealing your chance at glory

OpenStudy (anonymous):

which glory ?

OpenStudy (anonymous):

1 +2 + 4 + 8 +16 +....=-1

OpenStudy (anonymous):

Somming DIVERGENT SERIES

OpenStudy (anonymous):

thats not real :)

OpenStudy (anonymous):

Them Catalan numbers again..........:-)

OpenStudy (anonymous):

Neither are imaginary numbers......

OpenStudy (anonymous):

Yep the Catalan numbers are back @estudier

OpenStudy (anonymous):

@sauravshakya please read and apply to these two series

OpenStudy (anonymous):

or Mrs. @satellite73 or @experimentX ?

OpenStudy (experimentx):

lol ... you should be putting Mrs.

OpenStudy (anonymous):

second one is in the link you sent

OpenStudy (anonymous):

I meant- gentlemen ?

OpenStudy (anonymous):

a quick google search turns up euler mascheroni constant for harmonic series, which also seem so be \[\lim_{n\to \infty}\sum_{k=1}^nk-\ln(n)\]

OpenStudy (experimentx):

\[ \sum_{k=1}^\infty (-1)^{k+1} = 1 - 1 +1 - 1 + 1 ... \\ \] some guy had me shown some trick that it converges to 1/2 ... don't remember it though.

OpenStudy (anonymous):

that is NOT A SOLUTION!

OpenStudy (anonymous):

As you know \[ \ln(n) \to \infty\]

OpenStudy (anonymous):

So Mascheroni is NOT the answer

OpenStudy (experimentx):

\[ \text {Let } 1 - 1 +1 - 1 + 1 -1 + ... = a \\ 1 - (1 - 1 + 1 - 1 ...) = 1 - a = a \implies 2a = 1 \implies a = {1 \over 2 }\]

OpenStudy (experimentx):

this is false ... . of course :D

OpenStudy (experimentx):

By simple trick \[ \sum_{k=1}^\infty \frac{1}{k} \] it can be shown to be divergent ... for I don't remember any sleight of hand on this. Let's think for a while.

OpenStudy (anonymous):

FOR Ramanujan's sake READ !

OpenStudy (anonymous):

Messrs. @experimentX @satellite73 @sauravshakya

OpenStudy (anonymous):

1-1+2-6+24-120+......

OpenStudy (anonymous):

That will only work with Abel x> 5

OpenStudy (anonymous):

Too divergent - works with Double-Abel

OpenStudy (anonymous):

1-2+3-4+....

OpenStudy (anonymous):

Cesaro is enough for that

OpenStudy (anonymous):

The nasty one is 0.6 or thereabouts.....

OpenStudy (anonymous):

The nasty one is 0.6 or thereabouts.....

OpenStudy (anonymous):

Hey people - how can one cause you to read ?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Borel_summation

OpenStudy (experimentx):

Ah ... i was working on project euler problems.

OpenStudy (experimentx):

Man I flunked in life for the first time on on analysis.

OpenStudy (anonymous):

You need divergent sum therapy....:-)

OpenStudy (experimentx):

How do you plan to sum with Abel summation?

OpenStudy (anonymous):

\[\lim_{x\rightarrow 1-0} \sum_{k=0}^\infty a_k x^k = S.\]

OpenStudy (anonymous):

\[\Huge \lim_{x\rightarrow 1-0} \sum_{k=0}^\infty a_k x^k = S.\]

OpenStudy (experimentx):

go on ...

OpenStudy (anonymous):

."..it is no more doubtful that the sum of this series 1−2+3−4+5 + etc. is 1⁄4; since it arises from the expansion of the formula 1⁄(1+1)2, whose value is incontestably 1⁄4." Euler

OpenStudy (experimentx):

interesting ... how do you do that?

OpenStudy (experimentx):

1⁄(1+1)2 = 1/2 ( 1 - 1 + 1 - 1 + 1 - 1 ... )

OpenStudy (anonymous):

half of half is a quarter......!

OpenStudy (experimentx):

how did you manage to get n's in coefficient.

OpenStudy (experimentx):

let's tyr this \[ {1 \over (1 + x)^2} = \] http://www.wolframalpha.com/input/?i=expand+1%2F%281%2Bx%29^2 Idiot me!!

OpenStudy (experimentx):

should have noted.

OpenStudy (anonymous):

Scroll down to Euler and Borel

OpenStudy (anonymous):

Sorry, to Abel Summation....

OpenStudy (experimentx):

Interesting!!

OpenStudy (experimentx):

encountering this material for the first time man!!

OpenStudy (anonymous):

Good So just dive in and understand it "like a boss"...

OpenStudy (experimentx):

Hhaha i wish i could understand it "like a boss" ... I a bit slow. especially on algebra and analysis.

OpenStudy (anonymous):

Plug the sequence into the expression and try to find the limit

OpenStudy (anonymous):

The harmonic series is easily summed by Cesaro Summation

OpenStudy (anonymous):

Cesaro. Never heard of it. Thanks for bringing it accross.

OpenStudy (experimentx):

\[ \int {1 \over 1 -x}dx = \sum_{n=1}^\infty {x^n \over n }\] I wonder if this is useful.

OpenStudy (experimentx):

the partial sums are \[ {1}, {3 \over 4}, {11 \over 18}, {25 \over 48}, ...\]

OpenStudy (experimentx):

this seems to sonverge to zero as n->inf

OpenStudy (experimentx):

this is bugging me \[ \lim_{m\rightarrow \infty} {1 \over m }\sum_{n=1}^m {1 \over n } \sum_{k=1}^n {1 \over k}\]

OpenStudy (experimentx):

I wonder if i can use this http://en.wikipedia.org/wiki/Faulhaber%27s_formula